# HE-LRM: Encrypted Deep Learning Recommendation Models using Fully Homomorphic Encryption

### v1

@@ Abstract
Fully Homomorphic Encryption (FHE) is an encryption scheme that not only encrypts data but also allows for computations to be applied directly on the encrypted data. While computationally expensive, FHE can enable privacy-preserving neural inference in the client-server setting: a client encrypts their input with FHE and sends it to an untrusted server. The server then runs neural inference on the encrypted data and returns the encrypted results. The client decrypts the output locally, keeping both the input and result private from the server. Private inference has focused on networks with dense inputs such as image classification, and less attention has been given to networks with sparse features. Unlike dense inputs, sparse features require efficient encrypted lookup operations into large embedding tables, which present computational and memory constraints for FHE.
In this paper, we explore the challenges and opportunities when applying FHE to Deep Learning Recommendation Models (DLRM) from both a compiler and systems perspective. DLRMs utilize conventional MLPs for dense features and embedding tables to map sparse, categorical features to dense vector representations. We develop novel methods for performing compressed embedding lookups in order to reduce FHE computational costs while keeping the underlying model performant. Our embedding lookup improves upon a state-of-the-art approach by $\times7$ [1]. Furthermore, we present an efficient multi-embedding packing strategy that enables us to perform a 44 million parameter embedding lookup under FHE. Finally, we integrate our solutions into the open-source Orion framework and present HE-LRM, an end-to-end encrypted DLRM architecture. We evaluate HE-LRM on UCI (health prediction) and Criteo (click prediction), demonstrating that with the right compression and packing strategies, encrypted inference for recommendation systems is practical.
@@ 1 Introduction
Figure 1: Overview of Deep Learning Recommendation Models (DLRM) [2]. DLRMs process both dense and sparse features: dense features pass through an MLP (Bottom) whereas each sparse feature retrieves an embedding vector from a lookup table. The dense output and embeddings interact before passing through a final MLP (Top) to make a prediction. HE-LRM utilizes Fully Homomorphic Encryption (FHE) to perform end-to-end encrypted inference of this model, maintaining data privacy throughout the entire computation.
Fully Homomorphic Encryption (FHE) is a cryptographic scheme that allows computations to be performed directly on encrypted data without requiring any decryption. By enabling secure computations in untrusted environments, FHE offers a powerful solution for privacy-preserving applications in sensitive domains such as healthcare, finance, and personalized services. However, the high computational costs and the challenges of writing correct FHE programs limits its practicality to small-scale or narrowly scoped tasks.
Recent advances in FHE, from algorithmic improvements such as faster bootstrapping [3, 4, 5, 6, 7, 8], key-switching [9, 10], compiler optimizations [11, 12, 13, 14, 15, 16, 17, 18] and hardware acceleration [19, 20, 21, 22, 23, 24], have brought FHE closer to practical deployment for deep learning tasks. While prior works have explored FHE-based inference for simple neural networks or vision models, large-scale private recommendation models remain largely out of reach.
In particular, Deep Learning Recommendation Models (DLRMs), which enable personalized ranking and recommendation, pose unique challenges for privacy-preserving deployment. These models have the particularity that they process two types of inputs: dense numerical features (e.g., user age, device characteristics) and sparse categorical features (e.g., user location, user IDs, item categories), which index into large embedding tables. We refer to Figure 1 for an illustration of a general DLRM model. Embedding tables are often tied to sensitive user attributes such as demographics, behavior, and preferences. This motivates the need for a privacy-preserving DLRM capable of performing inference over encrypted data.
This paper presents a fully homomorphic encryption solution for DLRMs using the RNS-CKKS FHE scheme [25] to enable inference on encrypted user data without compromising privacy. To the best of our knowledge, no solution exists that runs such models under FHE. As mentioned above, a key challenge in deploying (private) DLRMs is that embedding tables can grow prohibitively large, and performing efficient indexing under FHE is difficult. To address this, we propose an improvement over Kim et al.’s solution [1] for homomorphic lookup table evaluation. We perform a similar compression of embedding tables but overcome many of their limitations such as wasted slots in a ciphertext, high CKKS multiplicative depth, and even security considerations, while at the same time allowing for a similar exponential compression factor. Furthermore, we extend our technique to include a multi-embedding packing strategy to operate over multiple embedding tables in a parallel fashion.
We design our FHE-based recommendation model, which we refer to as HE-LRM inspired by the DLRM architecture. We implement HE-LRM in the open-source Orion framework [18], leveraging their state-of-the-art packing strategies, automatic bootstrap placement, and scale management. Moreover, the Orion framework allows us to develop our HE-LRM components directly in PyTorch, allowing for faster iteration and rapid prototyping throughout the development process.
We demonstrate the efficiency of our solution by running HE-LRM on two datasets. First, we test our model on the (smaller) UCI health dataset to predict heart diseases. In this case, we use the $x^{2}$ activation function rather than ReLU. We report an FHE latency of approximately 24 seconds with a validation accuracy of 85%. We also run HE-LRM on the Criteo Kaggle dataset for which the size of the embedding tables becomes critically large. We define a threshold to determine which embedding tables to compress (when the number of rows exceeds the threshold) and varying our threshold, we demonstrate FHE latency from 488 seconds down to 227 seconds when the compression ratio is 31180×31180$\times$31180 ×, while the training loss and test AUC are minimally impacted by the compression. We also find the SiLU activation, which admits a lower-degree polynomial approximation than ReLU, significantly reduces the number of required bootstraps and improves efficiency. For example, the SiLU-based DLRM requires only 5 bootstraps and completes embedding lookups in 44 seconds, compared to 12 bootstraps and 51 seconds for its ReLU counterpart for a $10^{2}\times$ compressed model. Our main results are reported in Table I.
TABLE I: Summary of our experimental results when running HE-LRM on the UCI health dataset and the Criteo Kaggle dataset. Averaged over three runs.
[…]
$\mathsf{UCI}$ $\mathsf{Health}$ $224$ $24.22$
[…]
Finally, we provide a discussion of the broader limitations and challenges of private recommendation systems, highlighting both the promise and open questions in deploying such models securely. Concretely, we make the following contributions.
@@
    HE-LRM, an end-to-end FHE-based DLRM model that performs encrypted inference on both dense and sparse features.
@@
    An FHE-friendly embedding compression technique to efficiently achieve an exponential compression factor for large embedding tables that outperforms state-of-the-art by $\times7$.
@@
    A multi-embedding packing strategy that leverages a block-diagonal structure optimized for baby-step-giant-step linear transformations.

### v2

@@ Abstract
[…]
footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. Correspondence to: Anonymous Author <anon.email@domain.com>.
Preliminary work. Under review by the Machine Learning and Systems (MLSys) Conference. Do not distribute.
@@ 1 Introduction
Fully Homomorphic Encryption (FHE) is a cryptographic technique that enables computation directly on encrypted data without requiring decryption (10.1145/1536414.1536440). Given the extreme volume of neural network inferences processed daily (gupta2020architecturalimplicationsfacebooksdnnbased), FHE offers a solution for protecting sensitive user data in the cloud. While recent work has demonstrated FHE-based inference for CNNs and MLPs (gilad2016cryptonets; gazelle; orion), models with large embedding tables such as Deep Learning Recommendation Models (DLRMs) remain largely unexplored. As shown in Figure 1, DLRMs process sensitive user data such as demographic information, behavioral patterns, and personal preferences through embedding table lookups that transform these sparse categorical features into dense vectors (naumov2019deep).
However, a straightforward FHE approach requires encrypting massive one-hot vectors, resulting in prohibitively high computation and memory overhead. For example, the Criteo dataset contains embedding tables with over 33 million rows total, requiring thousands of FHE ciphertexts in order to perform embedding lookups for a single inference (criteo). This setup requires gigabytes of transmitted data, many expensive homomorphic operations, and latencies on the order of hours.
We develop a novel and high-performant FHE embedding table lookup optimization, which is an improvement over the prior state-of-the-art (kim2024privacypreserving). Our key insight is to develop an FHE-friendly embedding compression technique that maps original embedding tokens to deterministic coded tokens via digit decomposition that may be done client-side. We overcome several of their limitations such as underutilized ciphertext slots and a high multiplicative level consumption, while at the same time allowing for a similar exponential compression factor of embedding tables. We further propose a block-diagonal packing strategy to operate over multiple embedding tables in a parallel fashion. Our block-diagonal approach ensures that we maximize ciphertext slot utilization.
Using our FHE-friendly embedding compression technique, we present HE-LRM, a fully homomorphic solution for DLRMs to enable inference on encrypted user data without compromising privacy. We first validate our method on the UCI heart dataset and then apply our techniques to the industry-scale Criteo DLRM (heart_disease_45; criteo). While we target end-to-end encrypted DLRM inference, we note that embedding lookups are also required for transformer inference (transformer). Prior private transformer works bypass this challenge by performing embeddings client-side, exposing model parameters (nexus; thor). Our embedding solution is general and can be applied to these systems to keep embeddings server-side, improving their security.
Figure 1: Architecture of a Deep Learning Recommendation Model (DLRM). HE-LRM utilizes Fully Homomorphic Encryption (FHE) to perform end-to-end encrypted inference of this model, maintaining data privacy.
We make the following intellectual contributions:
  • HE-LRM: the first end-to-end FHE encrypted DLRM architecture supporting both dense and sparse features under FHE. We evaluate HE-LRM on two datasets, UCI Heart Disease and industry-scale Criteo.
  • An FHE-friendly embedding compression scheme that outperforms prior work kim2024privacypreserving by $77\times$ through client-side digit-decomposition. It consumes only one multiplicative level (vs. prior work’s $2+r+2s$) requiring fewer bootstraps, an expensive but necessary FHE operation for deep computations.
  • A block-diagonal embedding packing strategy to perform parallel lookups across multiple embedding tables, maximizing ciphertext slot utilization while remaining compatible with a state-of-the-art FHE linear transformation called baby-step giant-step (BSGS).
We evaluate HE-LRM on an industry-scale DLRM processing Criteo inputs, achieving end-to-end single-threaded CPU latencies ranging from 228 to 489 seconds depending upon embedding compression ratio. We also perform a case study projecting performance when FHE primitives are lowered to state-of-the-art GPU (arxiv-2024-cheddar) and custom ASIC implementations (ebel2024osiris). We find that hardware acceleration enables end-to-end latencies in the range of seconds to subseconds, making encrypted DLRM inference practical for real-world deployment.
2 Background
Figure 2: Single-threaded CPU latencies of primitive CKKS operations averaged over 30 runs. Both ciphertext-ciphertext (CT-CT) multiplication and ciphertext (CT) rotation require a compute and memory intensive key-switching operation. CKKS multiplies require ciphertexts to have at least two remaining levels.
This section provides background on the CKKS FHE scheme, the DLRM architecture, and the Orion framework for compiling PyTorch models to FHE.
2.1 Fully Homomorphic Encryption (CKKS)
We provide a systems-oriented overview of the CKKS FHE scheme. For details of the underlying cryptographic implementation, we refer the readers to the Appendix.
2.1.1 High-Level Overview
CKKS is an FHE encryption scheme that enables certain computations to be applied directly on encrypted vectors without requiring the secret key. Specifically, CKKS encrypts large floating-point vectors of a fixed power-of-two length (typically $n=2^{15}$ elements or slots) and enables SIMD addition, SIMD multiplication, and cyclic rotation upon these encrypted vectors. For addition and multiplication, one of the operands may be unencrypted (i.e., ciphertext-plaintext addition and multiplication), but the resulting output will be a ciphertext.
The atomic datatype in CKKS is a large vector of a fixed length; in other words, CKKS does not allow easy access to slicing or indexing of encrypted vectors. Despite these constraints, the CKKS instruction set (add, mul, and rotate) is sufficient to carry out complex operations such as linear layers and non-linear activation functions, making CKKS the de-facto FHE scheme for secure and outsourced deep learning applications.
2.1.2 Multiplicative Depth and Bootstrapping
Each ciphertext is associated with a multiplicative level, $\ell$, that determines the number of sequential multiplications it may undergo before decryption fails. In other words, the resulting output ciphertext of a CKKS multiplication has one less level than its inputs. When a ciphertext’s levels are exhausted, an expensive bootstrapping operation can restore levels to enable further computation. Bootstrapping is by far the most computationally expensive CKKS operation, taking roughly 20 seconds on a single-threaded CPU, and often dominates the latency of encrypted neural inference. For the cryptographic details on multiplicative depth and levels, we refer readers to Appendix A.
2.1.3 System-Level Costs
CKKS operations are both compute and memory intensive. Internally, ciphertexts are represented as integer polynomials with large coefficients (e.g., 100s to 1000s of bits) that are manipulated via modular arithmetic. As shown in Figure 2, all CKKS operations are substantially more expensive than their plaintext counterparts. Both ciphertext-ciphertext multiplication and ciphertext rotation are more expensive than all other CKKS operations due to a necessary but costly post-processing step called key-switching.
CKKS objects are also inherently large, with a single ciphertext typically around 10 to 30 MBs, and plaintext floating point vectors must be converted to plaintext CKKS objects, which also consume several MBs. Furthermore, the key-switching operation for ciphertext-ciphertext multiplication and ciphertext rotation require special evaluation keys ($\sim 200$ MB each), and a unique key is needed for each rotation offset. For this reason, most CKKS operations exceed server-grade last-level caches and require expensive DRAM accesses, further impacting performance.
2.1.4 Neural Network Layers in FHE
Linear layers such as fully connected or convolutional layers are implemented using the three CKKS primitives: SIMD addition, SIMD multiplication, and cyclic rotation. We refer the reader to Figure 2 of orion which illustrates two diagonal-based methods for performing a secure and outsourced linear transformation (i.e., a user sends their encrypted input to a server). The Halevi-Shoup method multiplies each diagonal of the plaintext weight matrix with the aligned input ciphertext, requiring $O(n)$ rotations. The Baby-Step Giant-Step (BSGS) optimization pre-rotates diagonals to reduce the rotation count to $O(\sqrt{n})$. Systems Implication: While linear layers only consume one multiplicative level, they require a large working memory for the evaluation keys for each rotation amount and for storing the plaintext weight matrix as its equivalent CKKS plaintext(s).
In contrast to linear layers, element-wise activation functions (e.g., ReLU or SiLU) must be approximated using high-degree polynomials that are comprised of SIMD multiplications and additions. The standard solution in CKKS is to use a composition of Chebyshev polynomials: for example, ReLU is commonly approximated using a composition of three polynomials of degrees 15, 15, and 27 lee2022. Systems Implication: Precise and accurate non-linear approximations necessitate high-degree polynomials which in turn have a high multiplicative depth. These approximations are the root cause of frequent bootstraps in FHE neural networks.
2.2 Deep Learning Recommendation Models
Deep Learning Recommendation Models (DLRMs) (naumov2019deep) are neural network architectures that accept as input both continuous (dense) and categorical (sparse) features and output probabilities for recommendation systems. Figure 1 shows an example of a typical DLRM: dense features are processed through an MLP whereas sparse features are transformed into dense vector representations via embedding table lookups. These processed inputs are then combined via an interaction operator after which they are passed through another MLP. The model outputs a probability corresponding to the likelihood of accepting a particular recommendation. In the context of advertising, the model predicts the likelihood of a user clicking an ad. More precisely, DLRMs consist of the following steps:
@@ 2.2 Deep Learning Recommendation Models
    The sparse input features are transformed into dense embedding vectors with each feature having a unique embedding table. For industry-scale DLRMs, embedding tables can contain millions of rows and consume GBs to TBs of memory (yin2021tt). These tables incur both large compute and memory costs in FHE, challenges which we address in this paper.